# Observable Propagation: A Data-Efficient Approach to Uncover Feature Vectors in Transformers

### v1

@@ Abstract
A key goal of current mechanistic interpretability research in NLP is to find linear features (also called “feature vectors”) for transformers: directions in activation space corresponding to concepts that are used by a given model in its computation. Present state-of-the-art methods for finding linear features require large amounts of labelled data – both laborious to acquire and computationally expensive to utilize. In this work, we introduce a novel method, called “observable propagation” (in short: ObProp), for finding linear features used by transformer language models in computing a given task – using almost no data. Our paradigm centers on the concept of “observables”, linear functionals corresponding to given tasks. We then introduce a mathematical theory for the analysis of feature vectors: we provide theoretical motivation for why LayerNorm nonlinearities do not affect the direction of feature vectors; we also introduce a similarity metric between feature vectors called the coupling coefficient which estimates the degree to which one feature’s output correlates with another’s. We use ObProp to perform extensive qualitative investigations into several tasks, including gendered occupational bias, political party prediction, and programming language detection. Our results suggest that ObProp surpasses traditional approaches for finding feature vectors in the low-data regime, and that ObProp can be used to better understand the mechanisms responsible for bias in large language models. Code for experiments can be found at github.com/jacobdunefsky/ObservablePropagation.
@@ 1 Introduction
Figure 1: Our method, ObProp, provides a paradigm for finding feature vectors used by a given transformer model for a given task. First, the task is translated into a linear functional on the logits of the model (part (a)); we call these linear functionals observables. Then, for a given computational path in the model, ObProp can find a feature vector used by that computational path to compute its output with respect to the observable (part (b)). A broader understanding of the computations that the model is carrying out can be obtained by comparing the feature vectors for different tasks. For instance, if the feature vectors for two different tasks have a high cosine similarity (part (c)), then this indicates that the model uses the same features for both tasks. We find that this is the case for the tasks of gendered pronoun prediction and occupation prediction, indicating that the model is biased, in that it uses gendered information to predict a person’s occupation.
@@
When a large language model predicts that the next token in a sentence is far more likely to be “him” than “her”, what is causing it to make this decision? The field of mechanistic interpretability aims to answer such questions by investigating how to decompose the computation carried out by a model into human-understandable pieces. This helps us predict their behavior, identify and correct discrepancies, align them with our goals, and assess their trustworthiness, especially in high-risk scenarios. The primary goal is to improve output prediction on real-world data distributions, identify and understand discrepancies between intended and actual behavior, align the model with our objectives, and assess trustworthiness in high-risk applications (Olah et al., 2018).
@@
One important notion in mechanistic interpretability is that of “features”. A feature can be thought of as a simple function of the activations at a particular layer of the model, the value of which is important for the model’s computation at that layer. For instance, in the textual domain, features used by a language model at some layer might reflect whether a token is an adverb, whether the language of the token is French, or other such characteristics. Possibly the most sought-after type of feature is a “linear feature”, or “feature vector”: a fixed vector in embedding space that the model utilizes by determining how much the input embedding points in the direction of that vector. Linear features are in some sense the holy grails of features: they are both easy for humans to interpret and amenable to mathematical analysis (Olah, 2022).
@@
Contributions    Our primary contribution is a method, which we call “observable propagation” (ObProp in short), for both finding feature vectors in large language models corresponding to given tasks, and analyzing these features in order to understand how they affect other tasks. Unlike non-feature-based interpretability methods such as saliency methods (Simonyan et al., 2013; Jacovi et al., 2021; Wallace et al., 2019) or circuit discovery methods (Conmy et al., 2023; Wang et al., 2022), observable propagation reveals the specific information from the model’s internal activations that are responsible for its output, rather than merely tokens or model components that are relevant. And unlike methods for finding feature vectors such as probing (Gurnee et al., 2023; Li et al., 2023; Elazar et al., 2021) or sparse autoencoders (Cunningham et al., 2023; Bricken et al., 2023), observable propagation can find these feature vectors without having to store large datasets of embeddings, perform many expensive forward passes, or utilize vast quantities of labeled data. In addition, we present the following contributions:
@@
  • We develop a detailed theoretical analysis of feature vectors. In Theorem 1, we provide theoretical motivation explaining why LayerNorm sublayers do not affect the direction of feature vectors; making progress towards answering the question of the extent to which LayerNorms are used in computation in transformers, which has been raised in mechanistic interpretability (Winsor, 2022). In Theorem 2, we introduce and motivate a measurement of feature vector similarity called the “coupling coefficient”, which can be used to determine the extent to which the model’s output on one task is coupled with the model’s output on another task.
@@ 1.1 Background and related work
In interpretability for NLP applications, there are a number of saliency-based methods that attempt to determine which tokens in the input are relevant to the model’s prediction (Simonyan et al., 2013; Jacovi et al., 2021; Wallace et al., 2019). Additionally, recent circuit-based mechanistic interpretability research has involved determining which components of a model are most relevant to the model’s computation on a given task (Conmy et al., 2023; Wang et al., 2022). Our work goes beyond these two approaches by considering not just relevant tokens, and not just relevant model components, but relevant feature vectors, which can be analyzed and compared to understand all the intermediate information used by models in their computation. A separate line of research aims to find feature vectors by performing supervised training of probes to find directions in embedding space that correspond to labels (Gurnee et al., 2023; Li et al., 2023; Elazar et al., 2021), or use unsupervised autoencoders on model embeddings to find feature vectors Cunningham et al. (2023); Bricken et al. (2023). ObProp does not rely on any training. A number of recent studies in interpretability involve finding feature vectors by decomposing transformer weight matrices into a set of basis vectors and projecting these vectors into token space (Dar et al., 2023; Millidge & Black, 2023). ObProp goes beyond this by taking into account nonlinearities, by finding precise feature vectors for tasks and by formulating the concept of “observables”, which is more general than the tasks considered in these prior works.
@@ 2 Observable propagation: from tasks to feature vectors
In this section, we present our method, which we call “observable propagation” (ObProp), for finding feature vectors directly corresponding to a given task. We begin by introducing the concept of “observables”, which is central to our paradigm. We then explain observable propagation for simple cases, and then build up to a general understanding.
@@ 2.1 Our paradigm: observables
Often, in mechanistic interpretability, we care about interpreting the model’s computation on a specific task. In particular, the model’s behavior on a task can frequently be expressed as the difference between the logits of two tokens. For instance, Mathwin et al. (2023) attempt to interpret the model’s understanding of gendered pronouns, and as such, measure the difference between the logits for the tokens " she" and " he". This has been identified as a general pattern of taking “logit differences” that appears in mechanistic interpretability work (Nanda, 2022).
@@
We thus define an observable to be a linear functional on the logits of a language model. In doing so, we no longer consider logit differences as merely a part of the process of performing an interpretability experiment; rather, we consider the broader class of linear functionals as being objects amenable to study in their own right. As we will see, concretizing observables like this will enable us to find sets of feature vectors corresponding to different observables.
@@ 2.2 Observable propagation for attention sublayers
Next, let us consider how to extend this idea to address attention sublayers in transformers. Attention combines information across tokens. They can be decomposed into two parts: the part that determines from which tokens information is taken (query-key interaction), and the part that determines what information is taken from each token to form the output (output-value). Elhage et al. (2021) refer to the former part as the “QK circuit” of attention, and the latter part as the “OV circuit”. Following their formulation, each attention layer can be written as
@@ 2.3 General form: addressing MLPs and LayerNorms
Along with attention sublayers, transformers also contain nonlinear MLP sublayers and LayerNorm nonlinearities before each sublayer. One main challenge in interpretability for large models has been the difficulty in understanding the MLP sublayers, due to the polysemantic nature of their neurons (Olah et al., 2020; Elhage et al., 2022). One prior approach to address this is modifying model architecture to increase the interpretability of MLP neurons (Elhage et al., 2022). Instead of architecture modification, we address these nonlinearities by approximating them as linear functions using their first-order Taylor approximations. This approach is reminiscent of that presented by Nanda et al. (2023), who use linearizations of language models to speed up the process of activation patching (Wang et al., 2022); we go beyond this by recognizing that the gradients used in these linearizations act as feature vectors that can be independently studied and interpreted, rather than merely making activation patching more efficient. Taking this into account, the general form of observable propagation, including first-order approximations of nonlinearities, can be implemented as follows. Consider a computational path $\mathcal{P}$ in the model through sublayers $l_{1}<l_{2}<\dots<l_{k}$. Then for a given observable $n$, the feature vector corresponding to sublayer $l$ in $\mathcal{P}$ can be computed according to Algorithm 1. Note that before every sublayer, there is a nonlinear LayerNorm operation. For greatest accuracy, one can find the feature vector corresponding to this LayerNorm by taking its gradient as described above. But as shown in Theorem 1, if one only cares about the directions of the feature vectors and not their magnitudes, then the LayerNorms can be ignored entirely.
@@
Let $W_{U}$ be the model unemebdding matrix.
if there exists a LayerNorm operation $x\mapsto f(x)$ before the unembedding operation then
     $y\leftarrow\nabla\left(n^{T}W_{U}f(x)\right)|_{x=x_{0}}$ for some suitable value of $x_{0}$
else
     $y\leftarrow(W_{U})^{T}n$
for $k\in\{\left|\mathcal{P}\right|,\dots,1\}$, starting at the end do
     if $l_{k}$ is an attention head then
         Let $W_{k}$ be the OV matrix for $l_{k}$.
         $y\leftarrow W_{k}^{T}y$.
     if $l_{k}$ is an a nonlinearity that maps $x\mapsto f(x)$ then
         $y\leftarrow\nabla\left(y^{T}W_{U}f(x)\right)|_{x=x_{0}}$ for some suitable value of $x_{0}$
$y$
@@ 2.4 The effect of LayerNorms on feature vectors
LayerNorms are ubiquitous in Transformers, appearing before every MLP and attention sublayer, and before the final unembedding matrix. Therefore, it is worth investigating how they affect feature vectors; if LayerNorms are highly nonlinear, this would be problematic.
The core LayerNorm function can be defined as $\operatorname{LayerNorm}(x)=\frac{Px}{\|Px\|}$ where $P$ is the orthogonal projection onto the hyperplane orthogonal to $\vec{1}$, the vector of all ones (Brody et al., 2023). Nanda et al. (2023) provide intuition for why we should expect that in high-dimensional spaces, LayerNorm is approximately linear. But it can be shown that the gradient of $\operatorname{LayerNorm}(x)$ is inversely proportional to $\left\|Px\right\|$ (see Appendix C.1), so we cannot consider LayerNorm gradients to be constant for inputs of different norms. However, empirically, we found that LayerNorms had almost no impact on the direction of feature vectors (see Appendix I). The following statement, which we prove in Appendix H, provides theoretical underpinning for this behavior:
@@
Define $f(x;n)=n\cdot\operatorname{LayerNorm}(x)$. Define
@@
$\mathbb{E}\left[\theta(x;n)\right]<2\arccos\left(\sqrt{1-1/(d-1)}\right)$
Note that after every LayerNorm as defined above, the output is multiplied by a fixed scalar constant equal to $\sqrt{d}$ (where $d$ is the embedding diension), multiplied by a learned diagonal matrix, and added to a learned vector. Thus, the actual operation implemented is $\sqrt{d}W\operatorname{LayerNorm}(x)+b$, where $W$ is the learned matrix and $b$ is the learned vector. Now, $b$ does not affect the gradient. Additionally, empirically, most of the entries in $W$ tend to be very close to one another (see Appendix E.2), which suggests that we can approximate $W$ as a scalar, meaning that $W$ primarily scales the gradient, rather than changing its direction. Therefore, if we want to analyze the directions of feature vectors rather than their magnitudes, then we can do so without worrying about LayerNorms.
@@ 3 Data-free analysis of feature vectors
An important question that we might want to ask about observables is the following: to what extent should we expect inputs that yield high outputs under observable $n_{1}$ to also yield high outputs for another observable $n_{2}$? If the outputs under $n_{1}$ and $n_{2}$ are highly correlated, then this suggests that the model uses the same underlying features for both observables. Having motivated this problem, let us translate it into the language of feature vectors. If $n_{1}$ and $n_{2}$ are observables with feature vectors $y_{1}$ and $y_{2}$ for a function $f$, then for inputs $x$, we have $n_{1}\cdot f(x)=y_{1}\cdot x$ and $n_{2}\cdot f(x)=y_{2}\cdot x$. Now, if we constrain our input $x$ to have norm $c$, and constrain $x\cdot y_{1}=k$, then what is the expected value of $x\cdot y_{2}$? And what are the maximum/minimum values of $x\cdot y_{2}$? We present the following theorem to answer both questions:
@@
We denote the value $\frac{y_{1}\cdot y_{2}}{\|y_{1}\|^{2}}$ by $C(y_{1},y_{2})$, and call it the “coupling coefficient from $y_{1}$ to $y_{2}$. Intuitively, $C(y_{1},y_{2})$ measures the expected dot product between a vector and $y_{2}$, assuming that that vector has a dot product of $1$ with $y_{1}$. Additionally, note that Theorem 2 also implies that the coupling coefficient becomes a more accurate estimator as the cosine similarity between $y_{1}$ and $y_{2}$ increases. This is borne out experimentally; see §4.1.

### v2

# Observable Propagation: Uncovering Feature Vectors in Transformers

@@ Abstract
A key goal of current mechanistic interpretability research in NLP is to find linear features (also called ?feature vectors?) for transformers: directions in activation space corresponding to concepts that are used by a given model in its computation. Present state-of-the-art methods for finding linear features require large amounts of labelled data – both laborious to acquire and computationally expensive to utilize. In this work, we introduce a novel method, called ?observable propagation? (in short: ObProp), for finding linear features used by transformer language models in computing a given task – using almost no data. Our paradigm centers on the concept of ?observables?, linear functionals corresponding to given tasks. We then introduce a mathematical theory for the analysis of feature vectors, including a similarity metric between feature vectors called the coupling coefficient which estimates the degree to which one feature’s output correlates with another’s. We use ObProp to perform extensive qualitative investigations into several tasks, including gendered occupational bias, political party prediction, and programming language detection. Our results suggest that ObProp surpasses traditional approaches for finding feature vectors in the low-data regime, and that ObProp can be used to better understand the mechanisms responsible for bias in large language models.
Interpretability
@@ 1 Introduction
When a large language model predicts that the next token in a sentence is far more likely to be ?him? than ?her?, what is causing it to make this decision? The field of mechanistic interpretability aims to answer such questions by investigating how to decompose the computation carried out by a model into human-understandable pieces. This helps us predict their behavior, identify and correct discrepancies, align them with our goals, and assess their trustworthiness, especially in high-risk scenarios. The primary goal is to improve output prediction on real-world data distributions, identify and understand discrepancies between intended and actual behavior, align the model with our objectives, and assess trustworthiness in high-risk applications (Olah et al., 2018).
@@
One important notion in mechanistic interpretability is that of ?features?. A feature can be thought of as a simple function of the activations at a particular layer of the model, the value of which is important for the model’s computation at that layer. For instance, in the textual domain, features used by a language model at some layer might reflect whether a token is an adverb, whether the language of the token is French, or other such characteristics. Possibly the most sought-after type of feature is a ?linear feature?, or ?feature vector?: a fixed vector in embedding space that the model utilizes by determining how much the input embedding points in the direction of that vector. Linear features are in some sense the holy grails of features: they are both easy for humans to interpret and amenable to mathematical analysis (Olah, 2022).
@@
Contributions    Our primary contribution is a method, which we call ?observable propagation? (ObProp in short), for both finding feature vectors in large language models corresponding to given tasks, and analyzing these features in order to understand how they affect other tasks. Unlike non-feature-based interpretability methods such as saliency methods (Simonyan et al., 2013; Jacovi et al., 2021; Wallace et al., 2019) or circuit discovery methods (Conmy et al., 2023; Wang et al., 2022), observable propagation reveals the specific information from the model’s internal activations that are responsible for its output, rather than merely tokens or model components that are relevant. And unlike methods for finding feature vectors such as probing (Gurnee et al., 2023; Li et al., 2023; Elazar et al., 2021) or sparse autoencoders (Cunningham et al., 2023), observable propagation can find these feature vectors without having to store large datasets of embeddings, perform many expensive forward passes, or utilize vast quantities of labeled data. In addition, we present the following contributions:
@@
  • We develop a detailed theoretical analysis of feature vectors. In Theorem 1, we provide theoretical motivation explaining why LayerNorm sublayers do not affect the direction of feature vectors; making progress towards answering the question of the extent to which LayerNorms are used in computation in transformers, which has been raised in mechanistic interpretability (Winsor, 2022). In Theorem 2, we introduce and motivate a measurement of feature vector similarity called the ?coupling coefficient?, which can be used to determine the extent to which the model’s output on one task is coupled with the model’s output on another task.
@@
All code used in this paper is provided at https://github.com/jacobdunefsky/ObservablePropagation.
@@ 1.1 Background and Related Work
In interpretability for NLP applications, there are a number of saliency-based methods that attempt to determine which tokens in the input are relevant to the model’s prediction (Simonyan et al., 2013; Jacovi et al., 2021; Wallace et al., 2019). Additionally, recent circuit-based mechanistic interpretability research has involved determining which components of a model are most relevant to the model’s computation on a given task (Conmy et al., 2023; Wang et al., 2022; Yu et al., 2023). Our work goes beyond these two approaches by considering not just relevant tokens, and not just relevant model components, but relevant feature vectors, which can be analyzed and compared to understand all the intermediate information used by models in their computation.
A separate line of research aims to find feature vectors by performing supervised training of probes to find directions in embedding space that correspond to labels (Gurnee et al., 2023; Li et al., 2023; Elazar et al., 2021; Tigges et al., 2023), or use unsupervised autoencoders on model embeddings to find feature vectors (Bricken et al., 2023; Cunningham et al., 2023). ObProp does not rely on any training – and importantly, exhibits greater fidelity to the model’s actual computation, because it directly uses the model weights to find feature vectors.
A number of recent studies in interpretability involve finding feature vectors by decomposing transformer weight matrices into a set of basis vectors and projecting these vectors into token space (Dar et al., 2023; Millidge & Black, 2023). ObProp goes beyond this by taking into account nonlinearities, by finding precise feature vectors for tasks (rather than merely being limited to choosing from among a fixed set of vectors), and by formulating the concept of ?observables?, which is more general than the tasks considered in these prior works.
Another approach to mechanistic interpretability involves making causal interventions on the model to determine whether a given abstraction accurately characterizes model behavior (Lepori et al., 2023; Wu et al., 2024). In contrast, ObProp directly finds specific feature vectors responsible for model behavior, rather than requiring humans to impose a given ontology onto the model to be tested.
Recently, Hernandez et al. (2023) used differentiation to obtain linear approximations to computations in language models responsible for encoding relations (such as ?plays musical instrument?) between entities. ObProp also uses differentiation to obtain linear approximations, but in contrast with this recent work, we seek to find feature vectors (rather than linear transformations) which are shared across multiple tasks rather than merely means by which a single task is implemented.
The concurrent work of Park et al. (2023) also investigates linear functionals on language models’ logit vectors and the connection between them and vectors in the model’s embedding space. However, their approach relies on sampling from the data distribution in order to estimate the covariance of unembedding vectors; this is in contrast to the data-free approach presented here.
@@ 2 Observable Propagation: From Tasks to Feature Vectors
In this section, we present our method, which we call ?observable propagation? (ObProp), for finding feature vectors directly corresponding to a given task. We begin by introducing the concept of ?observables?, which is central to our paradigm. We then explain observable propagation for simple cases, and then build up to a general understanding.
@@ 2.1 Our Paradigm: Observables
Often, in mechanistic interpretability, we care about interpreting the model’s computation on a specific task. In particular, the model’s behavior on a task can frequently be expressed as the difference between the logits of two tokens. For instance, (Mathwin et al., 2023) attempt to interpret the model’s understanding of gendered pronouns, and as such, measure the difference between the logits for the tokens " she" and " he". This has been identified as a general pattern of taking ?logit differences? that appears in mechanistic interpretability work (Nanda, 2022).
@@
We thus define an observable to be a linear functional on the logits of a language model. More formally:
Definition 2.1.
An observable is a linear functional [^1] Note that because all observables are linear functionals on a finite vector space, they can be written as row vectors. As such, it is often convenient to abuse notation, and associate an observable with its corresponding vector. $n:\mathbb{R}^{\mathtt{d\_vocab}}\to\mathbb{R}$, where d_vocab is the number of tokens in the model’s vocabulary. We refer to the action of taking the inner product of the model’s output logits with an observable $n$ as getting the output of the model under the observable $n$.
Returning to the previous example, the output of the model under the observable $(e_{\text{{" she"}}}-e_{\text{{" he"}}})$ corresponds to the logit difference between the " she" token and the " he" token.
In defining observables in this way, we no longer consider logit differences as merely a part of the process of performing an interpretability experiment; rather, we consider the broader class of linear functionals as being objects amenable to study in their own right. (And as we will see in §4.2, these linear functionals are not limited to merely those corresponding to logit differences between two tokens.) Next, we will demonstrate how concretizing observables like this enables us to find sets of feature vectors corresponding to different observables.
@@ 2.2 Observable Propagation for Attention Sublayers
Next, let us consider how to extend this idea to address attention sublayers in transformers. Attention combines information across tokens. They can be decomposed into two parts: the part that determines from which tokens information is taken (query-key interaction), and the part that determines what information is taken from each token to form the output (output-value). Elhage et al. (2021) refer to the former part as the ?QK circuit? of attention, and the latter part as the ?OV circuit?. Following their formulation, each attention layer can be written as
@@ 2.3 General Form: Addressing MLPs and LayerNorms
Along with attention sublayers, transformers also contain nonlinear MLP sublayers and LayerNorm nonlinearities before each sublayer. One main challenge in interpretability for large models has been the difficulty in understanding the MLP sublayers, due to the polysemantic nature of their neurons (Olah et al., 2020; Elhage et al., 2022). One prior approach to address this is modifying model architecture to increase the interpretability of MLP neurons (Elhage et al., 2022). Instead of architecture modification, we address these nonlinearities by approximating them as linear functions using their first-order Taylor approximations. This approach is reminiscent of that presented by Nanda et al. (2023), who use linearizations of language models to speed up the process of activation patching (Wang et al., 2022); we go beyond this by recognizing that the gradients used in these linearizations act as feature vectors that can be independently studied and interpreted, rather than merely making activation patching more efficient. Taking this into account, the general form of observable propagation, including first-order approximations of nonlinearities, can be implemented as follows. Consider a computational path $\mathcal{P}$ in the model through sublayers $l_{1}<l_{2}<\dots<l_{k}$. Then for a given observable $n$, the feature vector corresponding to sublayer $l$ in $\mathcal{P}$ can be computed according to Algorithm 1. (For details on how to choose $x_{0}$ in line 14, see App. C.)
Note that before every sublayer, there is a nonlinear LayerNorm operation. For greatest accuracy, one can find the feature vector corresponding to this LayerNorm by taking its gradient as described above. But as shown in Theorem 1, if one only cares about the directions of the feature vectors and not their magnitudes, then the LayerNorms can be ignored entirely.
@@
1:  Input: observable $n$
2:  Let $W_{U}$ be the model unemebdding matrix.
3:  if there exists a LayerNorm operation $x\mapsto f(x)$ before the unembedding operation then
4:    $y\leftarrow\nabla\left(n^{T}W_{U}f(x)\right)|_{x=x_{0}}$ for some suitable value of $x_{0}$
5:  else
6:    $y\leftarrow(W_{U})^{T}n$
7:  end if
8:  for $k\in\{\left|\mathcal{P}\right|,\dots,1\}$, starting at the end do
9:    if $l_{k}$ is an attention head then
10:       Let $W_{k}$ be the OV matrix for $l_{k}$.
11:       $y\leftarrow W_{k}^{T}y$.
12:    end if
13:    if $l_{k}$ is an a nonlinearity that maps $x\mapsto f(x)$ then
14:       $y\leftarrow\nabla\left(y^{T}f(x)\right)|_{x=x_{0}}$ for some suitable value of $x_{0}$
15:    end if
16:  end for
17:  Output: feature vector $y$
@@ 2.4 The Effect of LayerNorms on Feature Vectors
LayerNorm nonlinearities are ubiquitous in Transformers, appearing before every MLP and attention sublayer, and before the final unembedding matrix. Therefore, it is worth investigating how they affect feature vectors; if LayerNorms are highly nonlinear, this would cause trouble for ObProp.
Nanda et al. (2023) provide intuition for why we should expect that in high-dimensional spaces, LayerNorm is approximately linear. However, the gradient of LayerNorm depends on the norm of the input, so we cannot consider LayerNorm gradients to be constant for inputs of different norms. Nevertheless, empirically, we found that LayerNorms had almost no impact on the direction of feature vectors. In particular, for the feature vectors discussed in §4.3, we looked at the cosine similarities between the feature vectors computed by differentiating through LayerNorm and those computed by ignoring LayerNorm; the minimum cosine similarity was as high as 0.998. Further details can be found in Appendix E.6.
The following statement, which we prove in Appendix E.5, provides further theoretical underpinning for this behavior. (Note that while this theorem assumes that observables are normally distributed, which is not necessarily the case, we believe that it nevertheless provides useful theoretical motivation for explaining our empirical findings.)
@@
Define $\operatorname{LayerNorm}(x)=\frac{x-(\vec{1}^{T}x)\vec{1}}{\left\|x-(\vec{1}^{T}x)\vec{1}\right\|}$, where $\vec{1}$ is the vector of all ones. For a feature vector $n$, define $f(x;n)=n\cdot\operatorname{LayerNorm}(x)$. Define
@@
$\mathbb{E}\left[\theta(x;n)\right]<2\arccos\left(\sqrt{1-1/(d-1)}\right).$
@@ 3 Data-free Analysis of Feature Vectors
An important question that we might want to ask about a model’s behavior is the following: given two separate tasks, to what extent should we expect that the model will have a high output on one task whenever it has a high output on the other task? In other words, to what extent are the model’s outputs on the two tasks coupled? Let us translate this problem into the language of feature vectors. If $n_{1}$ and $n_{2}$ are observables with feature vectors $y_{1}$ and $y_{2}$ for a function $f$, then for inputs $x$, we have $n_{1}\cdot f(x)=y_{1}\cdot x$ and $n_{2}\cdot f(x)=y_{2}\cdot x$. Now, if we constrain our input $x$ to have norm $c$, and constrain $x\cdot y_{1}=k$, then what is the expected value of $x\cdot y_{2}$? And what are the maximum/minimum values of $x\cdot y_{2}$? We present the following theorem to provide theoretical grounding towards answering both questions:
@@
We denote the value $\frac{y_{1}\cdot y_{2}}{\|y_{1}\|^{2}}$ by $C(y_{1},y_{2})$, and call it the ?coupling coefficient from $y_{1}$ to $y_{2}$?. Intuitively, $C(y_{1},y_{2})$ measures the expected dot product between a vector and $y_{2}$, assuming that that vector has a dot product of $1$ with $y_{1}$. Additionally, note that Theorem 2 also implies that the coupling coefficient becomes a more accurate estimator as the cosine similarity between $y_{1}$ and $y_{2}$ increases.
Returning back to our original motivation, the coupling coefficient can be interpreted as estimating the constant of proportionality between a model’s outputs on two tasks (where each task corresponds to an observable and a feature vector); the cosine similarity can be interpreted as quantifying the extent to which the model’s outputs might deviate from this proportional relationship. In this manner, the coupling coefficient helps us predict the model’s behavior on unseen tasks.
Note that while Theorem 2 does make the assumption that hidden states $x$ are spherically-symmetrically distributed, nevertheless, experimental evidence does bear out that the coupling coefficient is an accurate estimator of the expected constant of proportionality between activations of different features, with accuracy increasing as the cosine similarity increases; see §4.1 for results.